OpenStudy (anonymous):

find the x and y intercepts for this equation -x – 3y = 9

hero (hero):

First, replace y with zero. Let me know what you get afterwards.

OpenStudy (anonymous):

-x-3(0)=9

hero (hero):

Very good. Now multiply -3 by zero. Let me know what you get afterwards.

OpenStudy (anonymous):

0

hero (hero):

Okay, but you should re-write the entire equation afterwards: -x - 0 = 9

hero (hero):

Next, factor out -1 from both terms on the left side. Let me know what you get.

OpenStudy (anonymous):

o

hero (hero):

When you factor out -1 on the left side you end up with -1(x + 0) = 9

hero (hero):

Now divide both sides by -1. Let me know what you get.

OpenStudy (anonymous):

-1/0=0 9/-1=-9

hero (hero):

You're dividing both sides by -1, which means the first fraction is not correct. Besides you can't divide by zero. It is illegal to do so.

OpenStudy (anonymous):

sorry I just dont think i am going to get this.

hero (hero):

It's not that difficult

hero (hero):

When you divide both sides by -1, you end up with x + 0 = -9

hero (hero):

Then the left side simplifies to just x = -9, which is your x-intercept.

hero (hero):

Now using the same original equation, set x = 0, then solve for y to find the y-intercept.

OpenStudy (anonymous):

0-3y=9

OpenStudy (anonymous):

or would it be 0+3y=9

mathslover (mathslover):

0-3y=9 go for this

OpenStudy (mertsj):

Find the x intercept by replacing y with 0. Find the y intercept by replacing x with 0.

OpenStudy (anonymous):

what would i do next

OpenStudy (mertsj):

Replace x with 0 and solve the equation for y. That will be the y intercept. Replace y with 0 and solve the equation for x. That will be the x intercept.

OpenStudy (mertsj):